If two charges of magnitudes 2c and 5c are located at free space between 10 metres with the constant k=20 metre/farad, then the force between the charges will be


1
Expert's answer
2021-01-28T12:19:25-0500

"F=k\frac{q_1q_2}{r^2}=9\cdot10^9\cdot\frac{2\cdot5}{10^2}=9\cdot10^{8}\ (N)"